OpenStudy (anonymous):

51 is the sum of 19 and jennelles savings : Translate into an equation

pooja195 (pooja195):

is means = sum means +

OpenStudy (anonymous):

51+19=j

OpenStudy (anonymous):

what does 51 mean?

pooja195 (pooja195):

leave it in the given order...

pooja195 (pooja195):

you have the info in the wrong order

OpenStudy (anonymous):

51 is basically your answer

OpenStudy (anonymous):

19+J=51

pooja195 (pooja195):

alex is says to write the expression not an answer.

OpenStudy (anonymous):

yup Cx

pooja195 (pooja195):

*it

OpenStudy (anonymous):

yeah i know that but 51 is the answer of the sum of 19 and j

pooja195 (pooja195):

jcolon its still wrong ..

OpenStudy (anonymous):

hes right

pooja195 (pooja195):

51 is the sum of 19 and j alex no he isnt.

OpenStudy (anonymous):

sum means whats added all together

pooja195 (pooja195):

im not dumb I know that but the expression the asker wrote is in the wrong order

OpenStudy (anonymous):

then whats the right order?

pooja195 (pooja195):

it says 51 is so 51=the sum of 19 and j alex im not a fan of giving the answer.I want the asker to learn.

OpenStudy (anonymous):

thats what he has

OpenStudy (anonymous):

look back

pooja195 (pooja195):

no he doesnt.

OpenStudy (anonymous):

he has 19 +J =51

pooja195 (pooja195):

thats wrong coz its flipped.

pooja195 (pooja195):

the order matters.

OpenStudy (anonymous):

i didnt know that matters

pooja195 (pooja195):

well now you know. :)